share out £80 between Tom and Jerry in the ratio 3:2
Answer by Seutip(231)

If they need to rationally distribute 80 to Tom and Jerry in ratio of 3:2, we need to do is
Let
3x - be Tom
2x - be Jerry
3x+2x=80
5x=80
x=16
3x = 3(16) Tom got 48
2x = 2(16) Jerry got 32
32+48 = 80
HOPE that helps!